# A Novel Terabit Grid-of-Beam Optical Wireless Multi-User Access Network With Beam Clustering

### v1

@@ IV-C2 Sum Rate Analysis
V Numerical Results and Discussions
[…]
V-A Clustering Scenarios: Spatial Distribution of Data Rate
In the baseline clustering scenario, every single beam is considered to be carrying independent data for its coverage spot. In this case, clusters are of size one, creating a pure space division multiple access (SDMA) network with the highest possible spatial reuse of the spectral resources. Since the bandwidth is fully reused across the network, the adjacent beams interfere with one another. Fig. LABEL:sub@Fig:6_NoClustering_a illustrates the approximate borders of the beam spots on the receiver plane. Fig. LABEL:sub@Fig:6_NoClustering_b demonstrates the spatial distribution of the received SINR on the same plane. The maximum SINR of $25$ dB is realized at the centers of the beam spots and the SINR value is on the order of $7$ dB at the edge of the network, i.e., in proximity to the walls. To obtain the spatial distribution of the achievable rate, the SINR values are mapped onto their respective data rates based on (52) by assuming that each cluster serves a single user [^2] Note that when there is a single user per cluster, achievable rates for NOMA and OFDMA are the same according to (52).. This hypothetical assumption helps to understand the performance bounds in different locations over the coverage area of the AP. According to Fig. LABEL:sub@Fig:6_NoClustering_c, data rates are about $10$ Gb/s to $15$ Gb/s at the centers of the beam spots except in locations near the walls where still a data rate of $8$ Gb/s is provided. For the baseline SDMA, the total number of clusters equals the total number of VCSELs in the AP, i.e., $N_{\mathrm{c}}=225$.
(a) Borders of SDMA clusters ($N_{\mathrm{c}}=225$)
(b) SINR
(c) Achievable rate
Figure 9: Baseline SDMA clustering and the resulting spatial distributions the received SINR and data rate on the receiver plane.
@@ V-A Clustering Scenarios: Spatial Distribution of Data Rate
Refer to caption (a) NOMA Refer to caption (b) OFDMA
Refer to caption (a) NOMA Refer to caption (b) OFDMA
Refer to caption (a) Average sum rate Refer to caption (b) Average fairness index
Figure 13: Overall network performance versus the user density $\rho$. Solid curves correspond to NOMA and dashed curves represent OFDMA.
@@ V-B Clustering Scenarios: Multi-User Performance Evaluation
We study the multi-user performance of the proposed clustering scenarios in terms of the sum rate and the fairness of the network based on Monte Carlo simulations by uniformly distributing the 2D coordinates of $K$ users in the room. The statistics of the performance metrics are evaluated over $10^{4}$ random realizations.
First, let us consider the case where the number of users is fixed at $K=10$ as a representative example of low number of users. The cumulative distribution functions of the sum rates offered by the clustering scenarios under NOMA and OFDMA are plotted in Fig. 13, where $\mathbb{P}$ denotes the probability of an event. In the case of NOMA, it can be observed from Fig. LABEL:sub@Fig:6_CDF1_10_SumRate_a that scenarios $1$ and $2$ result in a worse sum rate performance relative to the baseline SDMA, whereas scenarios $3$ and $4$ improve the performance. The sum rate for scenario $2$ closely follows that delivered by SDMA. However, for scenario $1$, there is a large gap as compared to SDMA. Unlike NOMA, when OFDMA is applied, as shown in Fig. LABEL:sub@Fig:6_CDF1_10_SumRate_b, it is evident that the sum rate for all four scenarios is always improved with respect to SDMA, as in OFDMA, the SINR improvement by CoMB-JT is a dominant factor leading to better data rates per user when the size of clusters $>1$.
@@
To further shed light into the multi-user performance, let us turn our attention to the average performance metrics. The average sum rate $\bar{R}$ is defined as the ergodic average of the sum rate, i.e., $\bar{R}=\mathbb{E}[R]$ where $\mathbb{E}$ denotes the expected value of a random variable. The average fairness index $\bar{J}$ is computed as $\bar{J}=\mathbb{E}[J]$. The results are mainly presented as a function of user density $\rho$, which is defined as the normalized number of users per unit area of the room:
@@
Fig. 13 shows the average performance metrics for both NOMA and OFDMA for comparison over a range of user densities up to $\rho=4$ $\text{user}/\text{m}^{2}$, equivalent to $K=100$ users. From Fig. LABEL:sub@Fig:6_AvgPerformance1_a, it can be observed that in both cases, the average sum rate is a monotonically increasing function of $\rho$. Also, NOMA clearly outperforms OFDMA in terms of the average sum rate for all the clustering scenarios. In the case of NOMA, for low user densities of $\rho\leq 0.4$ $\text{user}/\text{m}^{2}$, scenarios $3$ and $4$ slightly improve the performance over SDMA. By increasing $\rho$, SDMA takes the lead and its average sum rate approaches $1$ Tb/s for high user densities. Under OFDMA, all the scenarios provide gain with respect to SDMA over a range of $\rho\leq\rho_{\mathrm{th}}$, where $\rho_{\mathrm{th}}=0.6,1.8,2.1,3.6$ $\text{user}/\text{m}^{2}$ for scenarios $1$ to $4$, respectively. For both NOMA and OFDMA, as the user intensity increases, the average sum rate performance for all four scenarios converges to a constant value, which is scaled up with the number of clusters. Among the devised scenarios, scenario $4$ with $N_{\mathrm{c}}=49$ has the best performance and scenario $3$, though having the same number of clusters, holds the second best place. This highlights the point that the strategy used for adjusting the size of clusters in scenario $4$ is more efficient than that used in scenario $3$. We note that for both cases of NOMA and OFDMA, the sum rate is primarily limited in scenario $1$ in comparison with the other scenarios. The reason for this relates to the use of $9$ large clusters as shown in Fig. 10, which leads to a significant underutilisation of the spectral resources over the network. Also, since the channel gain is almost uniformly distributed over the area of each cluster in scenario $1$, the gain of NOMA over OFDMA for high user densities is less than when scenarios with smaller clusters are deployed.
From the perspective of the average fairness index as shown in Fig. LABEL:sub@Fig:6_AvgPerformance1_b, SDMA with NOMA retains the highest performance. For NOMA, all the scenarios exhibit a monotonically decreasing behavior with $\rho$. Under OFDMA, a similar trend is observed for all the scenarios excluding scenario 1, by which the fairness index raises for $\rho\geq 1.2$ $\text{user}/\text{m}^{2}$ and is notably greater than that achieved by the baseline SDMA at $\rho=4$ $\text{user}/\text{m}^{2}$. Furthermore, OFDMA performs better than NOMA in scenarios $1$ and $2$ and the performance gap of the two schemes becomes larger as $\rho$ increases. By comparison, NOMA brings about higher fairness indices in scenarios $3$ and $4$, and the performance gap between NOMA and OFDMA is consistently reduced when $\rho$ increases.
@@ VI Conclusions
We proposed the design of a novel double-tier AP architecture built upon a $3\times 3$ array of $5\times 5$ VCSEL arrays to establish an indoor grid-of-beam optical wireless access network. A full beam coverage with minimum overlaps between the beam spots were achieved by properly configuring the orientation angle of individual arrays. We introduced various beam clustering scenarios for downlink interference management and studied the multi-user performance of the network based on NOMA and OFDMA by considering a baseline SDMA scheme whereby the system bandwidth is reused for every single beam. The results evince the superior performance of SDMA with NOMA in terms of the sum rate and the fairness index especially for higher user densities. In this case, the average sum rate of the network approaches $1$ Tb/s when the user density exceeds $4$ $\text{user}/\text{m}^{2}$. Instead, performance gains of the clustering scenarios are evident for OFDMA. Particularly, scenario $4$ with $49$ clusters of balanced sizes yields $40\%$ improvement in the average sum rate performance compared to the baseline SDMA for a user density of $1.2$ $\text{user}/\text{m}^{2}$. On the other hand, scenario $1$ in which every VCSEL array is considered as a cluster of size $25$ improves the average fairness performance by $25\%$ relative to the baseline SDMA for a user density of $4$ $\text{user}/\text{m}^{2}$. These results are indicative of the effectiveness of clustering-based interference management in the OFDMA network. Future works include the optimization of the design parameters under quality of service constraints.

### v2

@@ V-A Clustering Scenarios: Spatial Distribution of Data Rate
(a) Average sum rate
(b) Average fairness index
Figure 13: Overall network performance versus the UE density $\rho$. Solid curves correspond to NOMA and dashed curves represent OFDMA.
@@ V-B Clustering Scenarios: Multi-User Performance Evaluation
We study the multi-user performance of the proposed clustering scenarios based on Monte Carlo simulations. To this end, the $xy$ coordinates of $K$ UEs are randomly distributed in the room with a uniform distribution, and the statistics of the sum rate and the fairness index of the network are evaluated by using (54) and (55) over $10^{4}$ random realizations. Note that the received SINR in (48) and (50) and hence the achievable rate in (53) primarily depend on the $xy$ coordinates of UEs through the channel gain $H_{kj}^{\mathcal{C}_{u}}$, which is given by (44), since $H_{kj}^{(i)}$ is related to the $xy$ coordinates of the $k$th UE via (42).
First, let us consider the case where the number of UEs is fixed at $K=10$ as a representative example of low number of UEs. The cumulative distribution functions of the sum rates offered by the clustering scenarios under NOMA and OFDMA are plotted in Fig. 13, where $\mathbb{P}$ denotes the probability of an event. In the case of NOMA, it can be observed from Fig. LABEL:sub@Fig:6_CDF1_10_SumRate_a that scenarios $1$ and $2$ result in a worse sum rate performance relative to the baseline SDMA, whereas scenarios $3$ and $4$ improve the performance. The sum rate for scenario $2$ closely follows that delivered by SDMA. However, for scenario $1$, there is a large gap as compared to SDMA. Unlike NOMA, when OFDMA is applied, as shown in Fig. LABEL:sub@Fig:6_CDF1_10_SumRate_b, it is evident that the sum rate for all four scenarios is always improved with respect to SDMA, as in OFDMA, the SINR improvement by CoMB-JT is a dominant factor leading to better data rates per UE when the size of clusters $>1$.
@@
To further shed light into the multi-user performance, let us turn our attention to the average performance metrics. The average sum rate $\bar{R}$ is defined as the ergodic average of the sum rate, i.e., $\bar{R}=\mathbb{E}[R]$ where $\mathbb{E}$ denotes the expected value of a random variable. The average fairness index $\bar{J}$ is computed as $\bar{J}=\mathbb{E}[J]$. The results are mainly presented as a function of UE density $\rho$, which is defined as the normalized number of UEs per unit area of the room:
@@
Fig. 13 shows the average performance metrics for both NOMA and OFDMA for comparison over a range of UE densities up to $\rho=4$ UE$/\text{m}^{2}$, equivalent to $K=100$ UEs. From Fig. LABEL:sub@Fig:6_AvgPerformance1_a, it can be observed that in both cases, the average sum rate is a monotonically increasing function of $\rho$. Also, NOMA clearly outperforms OFDMA in terms of the average sum rate for all the clustering scenarios. In the case of NOMA, for low UE densities of $\rho\leq 0.4$ UE$/\text{m}^{2}$, scenarios $3$ and $4$ slightly improve the performance over SDMA. By increasing $\rho$, SDMA takes the lead and its average sum rate approaches $1$ Tb/s for high UE densities. Under OFDMA, all the scenarios provide gain with respect to SDMA over a range of $\rho\leq\rho_{\mathrm{th}}$, where $\rho_{\mathrm{th}}=0.6,1.8,2.1,3.6$ UE$/\text{m}^{2}$ for scenarios $1$ to $4$, respectively. For both NOMA and OFDMA, as the UE density increases, the average sum rate performance for all four scenarios converges to a constant value, which is scaled up with the number of clusters. Among the devised scenarios, scenario $4$ with $N_{\mathrm{c}}=49$ has the best performance and scenario $3$, though having the same number of clusters, holds the second best place. This highlights the point that the strategy used for adjusting the size of clusters in scenario $4$ is more efficient than that used in scenario $3$. We note that for both cases of NOMA and OFDMA, the sum rate is primarily limited in scenario $1$ in comparison with the other scenarios. The reason for this relates to the use of $9$ large clusters as shown in Fig. 10, which leads to a significant underutilisation of the spectral resources over the network. Also, since the channel gain is almost uniformly distributed over the area of each cluster in scenario $1$, the gain of NOMA over OFDMA for high UE densities is less than when scenarios with smaller clusters are deployed.
From the perspective of the average fairness index as shown in Fig. LABEL:sub@Fig:6_AvgPerformance1_b, SDMA with NOMA retains the highest performance. For NOMA, all the scenarios exhibit a monotonically decreasing behavior with $\rho$. Under OFDMA, a similar trend is observed for all the scenarios excluding scenario 1, by which the fairness index raises for $\rho\geq 1.2$ UE$/\text{m}^{2}$ and is notably greater than that achieved by the baseline SDMA at $\rho=4$ UE$/\text{m}^{2}$. Furthermore, OFDMA performs better than NOMA in scenarios $1$ and $2$ and the performance gap of the two schemes becomes larger as $\rho$ increases. By comparison, NOMA brings about higher fairness indices in scenarios $3$ and $4$, and the performance gap between NOMA and OFDMA is consistently reduced when $\rho$ increases.
@@ VI Conclusions
We proposed the design of a novel double-tier AP architecture built upon a $3\times 3$ array of $5\times 5$ VCSEL arrays to establish an indoor grid-of-beam optical wireless access network. A full beam coverage with minimum overlaps between the beam spots were achieved by properly configuring the orientation angle of individual arrays. We introduced various beam clustering scenarios for downlink interference management and studied the multi-user performance of the network based on NOMA and OFDMA by considering a baseline SDMA scheme whereby the system bandwidth is reused for every single beam. The results evince the superior performance of SDMA with NOMA in terms of the sum rate and the fairness index especially for higher UE densities. In this case, the average sum rate of the network approaches $1$ Tb/s when the UE density exceeds $4$ UE$/\text{m}^{2}$. Instead, performance gains of the clustering scenarios are evident for OFDMA. Particularly, scenario $4$ with $49$ clusters of balanced sizes yields $40\%$ improvement in the average sum rate performance compared to the baseline SDMA for a UE density of $1.2$ UE$/\text{m}^{2}$. On the other hand, scenario $1$ in which every VCSEL array is considered as a cluster of size $25$ improves the average fairness performance by $25\%$ relative to the baseline SDMA for a UE density of $4$ UE$/\text{m}^{2}$. These results are indicative of the effectiveness of clustering-based interference management in the OFDMA network. Future works include investigating the impact of an imperfect knowledge of the UEs’ positions based on position estimation on the beam clustering performance.
Appendix A Derivation of the Refraction Vector in (10)
In order to derive (10), first, we establish a key identity that allows for the decomposition of any arbitrary vector $\mathbf{v}\in\mathbb{R}^{3}$ using a unit vector $\mathbf{n}\in\mathbb{R}^{3}$ in a given direction. Fig. LABEL:sub@Fig:VectorDecomposition depicts a right-handed Cartesian coordinate system with three unit and mutually orthogonal vectors of $\mathbf{r}$, $\mathbf{s}$ and $\mathbf{n}$, constituting an orthonormal basis for $\mathbb{R}^{3}$. Based on the right-hand rule, $\mathbf{s}=\mathbf{n}\times\mathbf{r}$ and $\mathbf{r}=\mathbf{s}\times\mathbf{n}$, and as a result:
$\mathbf{r}=\left(\mathbf{n}\times\mathbf{r}\right)\times\mathbf{n}.$ (57)
Note that the vector $\mathbf{v}$ lies in the plane spanned by $\mathbf{n}$ and $\mathbf{r}$, as represented by $P$ in Fig. LABEL:sub@Fig:VectorDecomposition. Therefore, it can be expressed in the form:
$\mathbf{v}=v_{\parallel}\mathbf{n}+v_{\perp}\mathbf{r},$ (58)
where $v_{\parallel}$ and $v_{\perp}$ are the components of $\mathbf{v}$ along the directions of $\mathbf{n}$ and $\mathbf{r}$, respectively. The projection of $\mathbf{v}$ onto the direction of $\mathbf{n}$ is given by:
$\mathbf{n}\cdot\mathbf{v}=v_{\parallel}\left(\mathbf{n}\cdot\mathbf{n}\right)+v_{\perp}\left(\mathbf{n}\cdot\mathbf{r}\right)=v_{\parallel},$ (59)
since $\mathbf{n}\cdot\mathbf{n}=1$ and $\mathbf{n}\cdot\mathbf{r}=0$. Also, the cross product of $\mathbf{n}$ and $\mathbf{v}$ is obtained as:
$\mathbf{n}\times\mathbf{v}=v_{\parallel}\left(\mathbf{n}\times\mathbf{n}\right)+v_{\perp}\left(\mathbf{n}\times\mathbf{r}\right)=v_{\perp}\left(\mathbf{n}\times\mathbf{r}\right),$ (60)
through the use of $\mathbf{n}\times\mathbf{n}=0$. Based on (57), it follows that:
$\left(\mathbf{n}\times\mathbf{v}\right)\times\mathbf{n}=v_{\perp}\left(\mathbf{n}\times\mathbf{r}\right)\times\mathbf{n}=v_{\perp}\mathbf{r}.$ (61)
Combining (60) and (61) with (58), we arrive at the following vector identity:
$\mathbf{v}=\left(\mathbf{n}\cdot\mathbf{v}\right)\mathbf{n}+\left(\mathbf{n}\times\mathbf{v}\right)\times\mathbf{n}.$ (62)
(a) Vector decomposition in $\mathbb{R}^{3}$
(b) The law of refraction
Figure 14: Vector representations used for the analysis: (a) The decomposition of a vector $\mathbf{v}$ using a unit vector $\mathbf{n}$ in a given direction; (b) The refraction of a light ray at the boundary surface $S$ of two homogeneous media of refractive indices $n_{1}$ and $n_{2}$ according to the law of refraction.
Now we can proceed with deriving the refraction vector $\mathbf{v}_{2}$ in terms of the incidence vector $\mathbf{v}_{1}$ and the normal vector $\mathbf{n}$ to the boundary surface $S$ between medium $1$ and medium $2$ with refractive indices $n_{1}$ and $n_{2}$, as shown in Fig. LABEL:sub@Fig:Law_of_Refraction. Note that $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ are unit vectors and they both lie on the plane of incidence $P$, forming the angle of incidence $\theta_{1}$ and the angle of refraction $\theta_{2}$ with respect to $\mathbf{n}$. By applying the vector decomposition in (62) and the law of refraction in (9), $\mathbf{v}_{2}$ can be expressed as:
$\mathbf{v}_{2}$ $=\left(\mathbf{n}\cdot\mathbf{v}_{2}\right)\mathbf{n}+\left(\mathbf{n}\times\mathbf{v}_{2}\right)\times\mathbf{n},$ (63)
$=\left(\mathbf{n}\cdot\mathbf{v}_{2}\right)\mathbf{n}+\mu\left(\mathbf{n}\times\mathbf{v}_{1}\right)\times\mathbf{n}.$
By using Snell’s law in (7) and considering that $0\leq\theta_{2}\leq\dfrac{\pi}{2}$ as observed from Fig. LABEL:sub@Fig:Law_of_Refraction, and hence $\cos\theta_{2}\geq 0$, we obtain:
$\mathbf{n}\cdot\mathbf{v}_{2}$ $=\cos\theta_{2},$ (64)
$=\sqrt{1-\sin^{2}\theta_{2}},$
$=\sqrt{1-\mu^{2}\sin^{2}\theta_{1}},$
$=\sqrt{1-\mu^{2}\left(1-\cos^{2}\theta_{1}\right)},$
$=\sqrt{1-\mu^{2}\left(1-\left(\mathbf{n}\cdot\mathbf{v}_{1}\right)^{2}\right)}.$
Furthermore, based on (62), we have:
$\left(\mathbf{n}\times\mathbf{v}_{1}\right)\times\mathbf{n}=\mathbf{v}_{1}-\mathbf{n}\cdot\mathbf{v}_{1}.$ (65)
Substituting (64) and (65) into (63), $\mathbf{v}_{2}$ is derived in terms of $\mathbf{n}$ and $\mathbf{v}_{1}$, resulting in the desired expression as given by (10):
$\mathbf{v}_{2}=\sqrt{1-\mu^{2}\left(1-\left(\mathbf{n}{\cdot}\mathbf{v}_{1}\right)^{2}\right)}\mathbf{n}+\mu\left(\mathbf{v}_{1}-\left(\mathbf{n}{\cdot}\mathbf{v}_{1}\right)\mathbf{n}\right).$ (66)